# Model Generalization on Text Attribute Graphs: Principles with Large Language Models

### v1

@@ 3.4 Principle II: Generalizable Graph Aggregation
Graph structures can provide essential correlated information for node label inference by characterizing the relationships between node attributes and labels [64, 65, 66, 67, 68, 69]. Specifically, we may consider each node’s label and attributes as random variables, and each edge as a coupling between them for connected node pairs. The fundamental BP algorithm enables principled statistical inference over this set of correlated random variables [39]. Since BP is inherently agnostic to the application domain of the TAG, emulating BP offers a mechanism to aggregate correlation information encoded in the graph structure across domains.
@@
Belief Propagation Exact inference for $\mathbb{P}_{\mathcal{G}}(Y|X)$ is intractable in large-scale graphs with cycles [70]. In practice, loopy belief propagation (LBP) is often used to conduct an approximate inference [39], which follows: Initialize the distributions $p_{j}^{(0)}(y_{j})\propto\varphi_{X_{i}}(y_{i})$ and $m_{i\to j}^{(0)}(y_{j})=1/c$ for all $i,j\in\mathcal{V}$. For $k=1,2,...,L$, we do
@@
where $\cong$ denotes the equality with difference up-to a constant. LSE stands for the log-sum-exp function: $\text{LSE}_{y_{j}}[f(y_{i},y_{j})]=\log\left[\sum_{y_{j}}\exp(f(y_{i},y_{j}))\right]$. The final $\arg\max_{y_{i}}p_{i}^{(k)}(y_{i})$ gives the label prediction. Detailed derivation can be found in Appendix. 8.
@@
[69] demonstrated that linear propagation can approximate a single iteration of LBP when feature quality is limited. Based on this insight, we adopt the following approximate LBP formulation (denoted as BP appr.):
@@ 4 Experiments
In this section, we evaluate LLM-BP based on its two design principles, with a primary focus on zero-shot node classification tasks. Evaluations of few-shot node classification and link prediction tasks are provided in Appendix. 9.4 9. First, we demonstrate the effectiveness of task-adaptive encoding and identify issues with existing methods that rely on aligning node embeddings with the LLM token space. Second, we validate the effectiveness of the proposed BP algorithm. Finally, we present the end-to-end performance of LLM-BP, comparing it to state-of-the-art baselines. We first introduce the datasets and baselines used in the study:
Datasets As listed in Table 1, we selected eleven real-world TAG datasets that encompass a variety of text domain shifts, including citation networks, e-commerce data, knowledge graphs, and webpage networks, which cover both homophily and heterophily structures. For more details, see Appendix 7.1.
@@
Cora [14] CS Publication Homopholic
Citeseer [15] CS Publication Homopholic
Pubmed [40] Medical Publication Homopholic
History [33] History Books Homopholic
Children [33] Children Literature Homopholic
Sportsfit [33] Sports Goods Homopholic
Wikics [34] Knowledge Graph Homopholic
Cornell [41] School Webpage Heterophilic
Texas [41] School Webpage Heterophilic
Wisconsin [41] School Webpage Heterophilic
Washington [41] School Webpage Heterophilic
@@
$\bullet$ Vanilla LM / LLM Encoders: including Sentence-BERT (SBert) [37], RoBERTa [38], text-embedding-3-large [71], and bge-en-icl [36], a state-of-the-art LLM2Vec encoder.
$\bullet$ Vanilla LLMs: including GPT-3.5-turbo [1] and GPT-4o [72], the latter being among the most advanced LLMs in reasoning. They process raw node texts without incorporating graph structures.
$\bullet$ Tuning LM Encoder / GNNs: including ZeroG [18], UniGLM [19] that tune LM encoders, ZeroG is specifically proposed for zero-shot node classification. DGI [21], GraphMAE [20] that perform Graph-SSL are also compared.
$\bullet$ LLMs with Graph Adapters: including LLaGA [26], TEA-GLM [29], and GraphGPT [27], which are the three representative works adopting LLMs with projectors to align compressed node representations with the token space.
$\bullet$ Multi-Task Graph Foundation Models: Consisting of OFA [55] and GOFA [57], which are the state-of-the-art multi-task foundation models.
$\bullet$ LLMs for Data Augmentation: referring to LLM-GNN [73], specifically designed for zero-shot node classification, which utilizes LLMs as annotators for pseudo-labels and further train GNNs for inference.
$\bullet$ Neighborhood Aggregation (NA): referring to the training-free method proposed in  [56], which injects graph structural information into node representations by directly aggregating the averaged neighborhood embeddings.
Settings: Unlike LLM-BP which is training-free, most of the baselines–except from vanilla encoders, LLMs or NA–require pre-training. Methods of vanilla encoders and LLM-BP that require sampling nodes to obtain class embeddings under zero-shot settings are repeated 30 times with seed 42 to 71, and the average performance is reported in the following experiment sections. Implementation details for baselines and LLM-BP can be found in Appendix. 7.3 7.2.
@@ 4.1 Evaluation for Task-Adaptive Node Embedding
$\bullet$ Exp.1: Ineffectiveness of LLMs w/ Graph Adapters Figure 3 illustrates the accuracy of encoder-based methods alongside two representative LLMs-with-graph-adapters methods across each dataset. Notably, using text embeddings generated by SBert [37] without incorporating graph structural information significantly outperforms both LLaGA [26] and GraphGPT [27]. These two methods align node representations that combine SBert embeddings with graph information to the LLMs’ token space via a projector. This finding suggests that the generalization capabilities of these approaches primarily stem from the pre-trained language model encoders rather than the LLMs’ inherent understanding of TAG data. Consequently, future works should exercise caution when adopting this strategy.
@@
According to Figure 3, the task-adaptive encoder achieves the best performance on most of the datasets, enhancing the vanilla LLM2Vec on average by $2.3\%$, highlighting the importance of incorporating task-specific information during encoding. To further illustrate this, we use the Citeseer [15] dataset as an example and perform t-SNE visualization [74] on the embeddings derived from the encoders. As shown in Fig. 3, when provided with class information, the task-adaptive encoder generates embeddings that exhibit tighter clustering for the same class compared to other baselines. The significance test of improvement from task-adatove encoding is provided in Table. 5 in Appendix. 5.
Note that the benefits of class information are observed only in encoders derived from LLM decoders potentially due to their strong contextual learning capabilities. As illustrated in Fig. 4, incorporating class information into smaller LM encoders, such as SBert [37] or RoBERTa [38], may even degrade performance. Regarding Text-embedding-3-large [71], the impact of class information remains inconclusive due to the unknown internal mechanisms of the black-box encoder.
@@ 4.2 Generalizable Graph Aggregation
Sbert [37] 69.75 67.21 66.69 63.31 70.57 71.38 53.53 20.45 22.59 20.13 43.79 38.26 59.06 56.19 63.66 54.39 64.58 49.79 62.10 52.07 63.52 48.00 7.27 7.09
Roberta [38] 70.71 68.47 66.95 63.57 69.54 70.31 55.39 21.84 24.25 22.41 41.51 36.09 59.08 56.49 61.68 51.84 62.25 49.26 60.33 49.08 60.60 45.34 7.18 7.18
Text-Embedding-3-Large [71] 71.90 69.87 66.24 63.30 75.96 75.75 50.15 19.21 24.68 24.10 58.39 53.03 61.78 58.82 81.50 70.11 75.42 63.17 73.14 63.02 66.35 57.69 5.36 4.45
LLM2Vec [35] 67.34 65.92 67.13 64.37 74.57 74.65 53.14 19.06 25.56 24.31 57.00 52.29 62.34 58.32 81.26 69.08 76.68 63.12 73.36 62.50 65.92 53.34 5.64 5.36
SBert + NA [56] 72.49 69.90 68.66 64.75 71.26 71.87 57.86 21.98 25.28 22.74 46.84 40.85 66.26 63.57 54.21 44.66 56.04 41.09 54.23 46.11 58.88 43.05 5.82 6.00
GPT-3.5-turbo [1] 70.11 52.11 66.83 47.58 89.75 66.16 55.07 30.36 29.73 26.13 67.21 54.45 65.53 51.19 45.54 39.30 56.14 32.53 58.86 46.84 51.09 35.68 5.64 8.18
GPT-4o [72] 70.29 62.95 64.77 47.78 89.85 67.39 53.30 31.68 30.76 29.20 66.35 56.22 66.10 56.04 45.54 41.92 63.10 50.51 56.60 52.54 48.90 42.54 5.91 6.36
UniGLM [19] 45.57 43.25 52.26 48.41 70.33 69.78 44.24 24.84 21.48 19.17 33.46 32.99 55.05 52.08 23.03 22.06 21.39 18.90 27.16 26.45 24.01 23.08 11.36 9.91
ZeroG [18] 60.4 56.02 50.35 45.15 74.68 71.75 36.55 16.84 12.72 12.61 14.27 5.33 46.74 40.86 10.47 6.46 53.48 15.95 12.66 5.02 8.3 3.07 12.27 12.73
DGI [21] 16.79 12.77 15.24 15.04 25.10 19.18 20.98 3.89 2.22 1.04 7.48 3.47 14.98 4.24 14.66 10.02 11.23 9.42 12.08 6.95 20.96 14.15 13.91 14.73
GraphMAE [20] 15.13 7.10 8.11 7.67 36.56 34.29 36.36 5.75 7.24 1.97 30.50 6.99 8.91 4.03 23.04 14.95 17.65 11.67 23.02 11.87 24.89 13.34 15.18 15.45
OFA [55] 20.36 16.57 41.31 33.37 28.18 26.62 8.25 3.48 3.05 2.29 15.18 4.7 30.77 25.22 29.84 12.62 11.77 5.87 4.8 3.44 6.04 4.28 13.91 14.73
GOFA [57] 71.06 70.21 65.72 64.18 74.76 73.00 56.25 31.57 12.15 7.73 37.87 33.19 68.62 62.93 39.50 35.47 38.37 29.54 32.51 25.12 31.02 21.24 8.00 7.45
GraphGPT [27] 17.48 12.68 13.93 12.78 42.94 25.68 12.31 9.15 9.94 4.24 4.53 2.44 33.59 30.21 10.18 14.71 18.48 9.85 12.35 6.32 20.64 15.79 14.55 14.64
LLAGA [26] 11.62 14.42 19.52 23.34 7.56 13.42 7.95 8.89 10.09 5.02 1.84 2.66 10.98 16.73 12.57 20.1 15.51 22.97 15.09 20.85 10.48 18.98 15.36 13.64
@@
Figure 6: Experiments on graph information aggregation. ‘Raw’ refers no graph structure usage, ‘w/ NA’ refers to the neighborhood embedding aggregation (NA) proposed in [56], ‘w/ BP’ refers to the belief propagation following Eq. 6, ‘w/ BP (appr.)’ refers to its simplified linear form that follows Eq. 7.
$\bullet$ Exp.3: LLM Agents for Homophily Level $r$ Estimation As shown in Fig.5 (Left), we randomly sample $k$ edges ($k=100$ for large graphs and $k=50$ for small ones), incorporating them into prompts (Fig.1) for LLM-based estimation of the homophily level $r$ (Sec.3.4). We evaluate four LLMs: GPT-4o, GPT-4o-mini[72], GPT-3.5-Turbo [1], and Mistral-7B-Instruct v0.3 [75]. Each model responds to each node pair over five trials, with the final estimate determined by majority voting. Full results are provided in Fig.7 in appendix, demonstrating that GPT-4o-mini and GPT-4o effectively estimate $r$, GPT-3.5-Turbo performs reasonably well, while Mistral-7B-Instruct-v0.3 fails. Balancing accuracy and cost efficiency, we select GPT-4o-mini’s estimation (Fig. 5 Right) for subsequent studies.
@@ 4.3 End-to-End Evaluation
$\bullet$ Exp.5: Main Results in the Zero-Shot Setting The main experimental results are presented in Table 2. Among the baselines, vanilla encoders and LLMs demonstrate strong zero-shot generalization. GPT-3.5-Turbo [1] ranks first on the Sportsfit dataset, while GPT-4o [72] achieves the best performance on Pubmed and Children.
UniGLM [19] and ZeroG [18] perform well in domains aligned with their pre-training, such as citation networks (e.g., ZeroG enhances SBert’s performance on Cora, Pubmed, and Wikics). However, both struggle on TAGs with unseen text distributions (e.g., Sportsfit) or novel graph structures (e.g., webpage networks), suggesting that fine-tuned LM encoders may suffer performance degradation on out-of-domain TAGs. Similarly, graph-SSL methods (DGI [21], GraphMAE [20]) show limited generalization across structural shifts.
Among multi-task graph foundation models, GOFA achieves strong performance, likely benefiting from a larger pre-training corpus for graph-text alignment [76, 77] compared to GraphGPT [27] and LLaGA [26], which are trained solely on ogbn-arxiv. However, GOFA still requires broader pre-training and instruction fine-tuning to improve generalization under text domain shifts, and its reliance on GNNs may limit effectiveness on heterophilic data.
Notably, LLM-BP and LLM-BP (appr.) achieve the highest average ranking across all datasets. Another interesting observation is that when we randomly sample $20c$ nodes to obtain the class embeddings with the help of LLMs following Algorithm. 1, the zero-shot performance of the encoders in this setting is comparable to their performance between $5$-$10$-shot setting as shown in Table. 8 in the Appendix. Further comparisons with LLM-GNN [73] and TEA-GLM [29] are provided in Appendix 9.3.
$\bullet$ Exp.6: Main Results under Few-shot Setting We conduct the evaluation in $k=1,3,5,10$-shot settings. Using 10 different random seeds, we sample the shots from the training set and repeat the experiments 10 times. The experimental results are presented in Table 8 in Appendix 9. Across all $k$-shot settings, LLM-BP and LLM-BP (appr.) outperform the baseline models.

### v2

@@ 3.4 Principle II: Generalizable Graph Aggregation
Graph structures can provide essential correlated information for node label inference by characterizing the relationships between node attributes and labels (Zhu et al., 2003; Kipf & Welling, 2016; Veličković et al., 2017; Hamilton et al., 2017; Zhu et al., 2020; Wei et al., 2022).
Specifically, we may consider each node’s label and attributes as random variables, and each edge as a coupling between them for connected node pairs. The fundamental BP algorithm enables principled statistical inference over this set of correlated random variables (Murphy et al., 2013). Since BP is inherently agnostic to the application domain of the TAG, emulating BP offers a mechanism to aggregate correlation information encoded in the graph structure across domains.
@@
Belief Propagation Exact inference for $\mathbb{P}_{\mathcal{G}}(Y|X)$ is intractable in large-scale graphs with cycles (Koller, 2009). In practice, loopy belief propagation (LBP) is often used to conduct an approximate inference (Murphy et al., 2013), which follows: Initialize the distributions $p_{j}^{(0)}(y_{j})\propto\varphi_{X_{i}}(y_{i})$ and $m_{i\to j}^{(0)}(y_{j})=1/c$ for all $i,j\in\mathcal{V}$. For $k=1,2,...,L$, we do
@@
where $\cong$ denotes the equality with difference up-to a constant. LSE stands for the log-sum-exp function: $\text{LSE}_{y_{j}}[f(y_{i},y_{j})]=\log\left[\sum_{y_{j}}\exp(f(y_{i},y_{j}))\right]$. The final $\arg\max_{y_{i}}p_{i}^{(k)}(y_{i})$ gives the label prediction. Detailed derivation can be found in Appendix. C.
@@
(Wei et al., 2022) demonstrated that linear propagation can approximate a single iteration of LBP when feature quality is limited. Based on this insight, we adopt the following approximate LBP formulation (denoted as BP appr.):
@@ 4 Experiments
In this section, we evaluate LLM-BP based on its two design principles, with a primary focus on zero-shot node classification tasks. Evaluations of few-shot node classification and link prediction tasks are provided in Appendix. D.6 11. First, we demonstrate the effectiveness of task-adaptive encoding and identify issues with existing methods that rely on aligning node embeddings with the LLM token space. Second, we validate the effectiveness of the proposed BP algorithm. Finally, we present the end-to-end performance of LLM-BP, comparing it to state-of-the-art baselines. We first introduce the datasets and baselines used in the study:
Datasets As listed in Table 1, we selected eleven real-world TAG datasets that encompass a variety of text domain shifts, including citation networks, e-commerce data, knowledge graphs, and webpage networks, which cover both homophily and heterophily structures. For more details, see Appendix B.1.
@@
Cora (McCallum et al., 2000) CS Publication Homopholic
Citeseer (Giles et al., 1998) CS Publication Homopholic
Pubmed (Sen et al., 2008) Medical Publication Homopholic
History (Ni et al., 2019) History Books Homopholic
Children (Ni et al., 2019) Children Literature Homopholic
Sportsfit (Ni et al., 2019) Sports Goods Homopholic
Wikics (Mernyei & Cangea, 2020) Knowledge Graph Homopholic
Cornell (Craven et al., 1998) School Webpage Heterophilic
Texas (Craven et al., 1998) School Webpage Heterophilic
Wisconsin (Craven et al., 1998) School Webpage Heterophilic
Washington (Craven et al., 1998) School Webpage Heterophilic
@@
$\bullet$ Vanilla LM / LLM Encoders: including Sentence-BERT (SBert) (Reimers, 2019), RoBERTa (Liu, 2019), text-embedding-3-large (OpenAI, 2024), and bge-en-icl (Li et al., 2024a), a state-of-the-art LLM2Vec encoder.
$\bullet$ Vanilla LLMs: including GPT-3.5-turbo (Achiam et al., 2023) and GPT-4o (Hurst et al., 2024), the latter being among the most advanced LLMs in reasoning. They process raw node texts without incorporating graph structures.
$\bullet$ Tuning LM Encoder / GNNs: including ZeroG (Li et al., 2024e), UniGLM (Fang et al., 2024) that tune LM encoders, ZeroG is specifically proposed for zero-shot node classification. DGI (Veličković et al., 2018), GraphMAE (Hou et al., 2022) that perform Graph-SSL are also compared.
$\bullet$ LLMs with Graph Adapters: including LLaGA (Chen et al., 2024b), TEA-GLM (Wang et al., 2024), and GraphGPT (Tang et al., 2024a), which are the three representative works adopting LLMs with projectors to align compressed node representations with the token space.
$\bullet$ Multi-Task Graph Foundation Models: Consisting of OFA (Liu et al., 2023a) and GOFA (Kong et al., 2024), which are the state-of-the-art multi-task foundation models.
$\bullet$ LLMs for Data Augmentation: referring to LLM-GNN (Chen et al., 2024e), specifically designed for zero-shot node classification, which utilizes LLMs as annotators for pseudo-labels and further train GNNs for inference.
$\bullet$ Neighborhood Aggregation (NA): referring to the training-free method proposed in  (Yang et al., 2024), which injects graph structural information into node representations by directly aggregating the averaged neighborhood embeddings.
Settings: Unlike LLM-BP which does not require additional fine-tuning of LLMs, most of the baselines–except from vanilla encoders, LLMs or NA–require fine-tuning. Methods of vanilla encoders and LLM-BP that require sampling nodes to obtain class embeddings under zero-shot settings are repeated 30 times with seed 42 to 71, and the average performance is reported in the following experiment sections. Implementation details for baselines and LLM-BP can be found in Appendix. B.3 B.2.
@@ 4.1 Evaluation for Task-Adaptive Node Embedding
$\bullet$ Exp.1: Ineffectiveness of LLMs w/ Graph Adapters
Figure 3 illustrates the accuracy of encoder-based methods alongside two representative LLMs-with-graph-adapters methods across each dataset. Notably, using text embeddings generated by SBert (Reimers, 2019) without incorporating graph structural information significantly outperforms both LLaGA (Chen et al., 2024b) and GraphGPT (Tang et al., 2024a). These two methods align node representations that combine SBert embeddings with graph information to the LLMs’ token space via a projector. This finding suggests that the generalization capabilities of these approaches primarily stem from the pre-trained language model encoders rather than the LLMs’ inherent understanding of TAG data. Consequently, future works should exercise caution when adopting this strategy.
@@
According to Figure 3, the task-adaptive encoder achieves the best performance on most of the datasets, enhancing the vanilla LLM2Vec on average by $2.3\%$, highlighting the importance of incorporating task-specific information during encoding. To further illustrate this, we use the Citeseer (Giles et al., 1998) dataset as an example and perform t-SNE visualization (Van der Maaten & Hinton, 2008) on the embeddings derived from the encoders. As shown in Fig. 3, when provided with class information, the task-adaptive encoder generates embeddings that exhibit tighter clustering for the same class compared to other baselines. The significance test of improvement from task-adatove encoding is provided in Table. 5 in Appendix. 5.
Note that the benefits of class information are observed only in encoders derived from LLM decoders potentially due to their strong contextual learning capabilities. As illustrated in Fig. 4, incorporating class information into smaller LM encoders, such as SBert (Reimers, 2019) or RoBERTa (Liu, 2019), may even degrade performance. Regarding Text-embedding-3-large (OpenAI, 2024), the impact of class information remains inconclusive due to the unknown internal mechanisms of the black-box encoder.
@@ 4.2 Generalizable Graph Aggregation
Sbert (Reimers, 2019) 69.75 67.21 66.69 63.31 70.57 71.38 53.53 20.45 22.59 20.13 43.79 38.26 59.06 56.19 63.66 54.39 64.58 49.79 62.10 52.07 63.52 48.00 7.27 7.09
Roberta (Liu, 2019) 70.71 68.47 66.95 63.57 69.54 70.31 55.39 21.84 24.25 22.41 41.51 36.09 59.08 56.49 61.68 51.84 62.25 49.26 60.33 49.08 60.60 45.34 7.18 7.18
Text-Embedding-3-Large (OpenAI, 2024) 71.90 69.87 66.24 63.30 75.96 75.75 50.15 19.21 24.68 24.10 58.39 53.03 61.78 58.82 81.50 70.11 75.42 63.17 73.14 63.02 66.35 57.69 5.36 4.45
LLM2Vec (BehnamGhader et al., 2024) 67.34 65.92 67.13 64.37 74.57 74.65 53.14 19.06 25.56 24.31 57.00 52.29 62.34 58.32 81.26 69.08 76.68 63.12 73.36 62.50 65.92 53.34 5.64 5.36
SBert + NA (Yang et al., 2024) 72.49 69.90 68.66 64.75 71.26 71.87 57.86 21.98 25.28 22.74 46.84 40.85 66.26 63.57 54.21 44.66 56.04 41.09 54.23 46.11 58.88 43.05 5.82 6.00
GPT-3.5-turbo (Achiam et al., 2023) 70.11 52.11 66.83 47.58 89.75 66.16 55.07 30.36 29.73 26.13 67.21 54.45 65.53 51.19 45.54 39.30 56.14 32.53 58.86 46.84 51.09 35.68 5.64 8.18
GPT-4o (Hurst et al., 2024) 70.29 62.95 64.77 47.78 89.85 67.39 53.30 31.68 30.76 29.20 66.35 56.22 66.10 56.04 45.54 41.92 63.10 50.51 56.60 52.54 48.90 42.54 5.91 6.36
UniGLM (Fang et al., 2024) 45.57 43.25 52.26 48.41 70.33 69.78 44.24 24.84 21.48 19.17 33.46 32.99 55.05 52.08 23.03 22.06 21.39 18.90 27.16 26.45 24.01 23.08 11.36 9.91
ZeroG (Li et al., 2024e) 60.4 56.02 50.35 45.15 74.68 71.75 36.55 16.84 12.72 12.61 14.27 5.33 46.74 40.86 10.47 6.46 53.48 15.95 12.66 5.02 8.3 3.07 12.27 12.73
DGI (Veličković et al., 2018) 16.79 12.77 15.24 15.04 25.10 19.18 20.98 3.89 2.22 1.04 7.48 3.47 14.98 4.24 14.66 10.02 11.23 9.42 12.08 6.95 20.96 14.15 13.91 14.73
GraphMAE (Hou et al., 2022) 15.13 7.10 8.11 7.67 36.56 34.29 36.36 5.75 7.24 1.97 30.50 6.99 8.91 4.03 23.04 14.95 17.65 11.67 23.02 11.87 24.89 13.34 15.18 15.45
OFA (Liu et al., 2023a) 20.36 16.57 41.31 33.37 28.18 26.62 8.25 3.48 3.05 2.29 15.18 4.7 30.77 25.22 29.84 12.62 11.77 5.87 4.8 3.44 6.04 4.28 13.91 14.73
GOFA (Kong et al., 2024) 71.06 70.21 65.72 64.18 74.76 73.00 56.25 31.57 12.15 7.73 37.87 33.19 68.62 62.93 39.50 35.47 38.37 29.54 32.51 25.12 31.02 21.24 8.00 7.45
GraphGPT (Tang et al., 2024a) 17.48 12.68 13.93 12.78 42.94 25.68 12.31 9.15 9.94 4.24 4.53 2.44 33.59 30.21 10.18 14.71 18.48 9.85 12.35 6.32 20.64 15.79 14.55 14.64
LLAGA (Chen et al., 2024b) 11.62 14.42 19.52 23.34 7.56 13.42 7.95 8.89 10.09 5.02 1.84 2.66 10.98 16.73 12.57 20.1 15.51 22.97 15.09 20.85 10.48 18.98 15.36 13.64
@@
Figure 6: Experiments on graph information aggregation. ‘Raw’ refers no graph structure usage, ‘w/ NA’ refers to the neighborhood embedding aggregation (NA) proposed in (Yang et al., 2024), ‘w/ BP’ refers to the belief propagation following Eq. 6, ‘w/ BP (appr.)’ refers to its simplified linear form that follows Eq. 7.
$\bullet$ Exp.3: LLM Agents for Homophily Level $r$ Estimation As shown in Fig.5 (Left), we randomly sample $k$ edges ($k=100$ for large graphs and $k=50$ for small ones), incorporating them into prompts (Fig.1) for LLM-based estimation of the homophily level $r$ (Sec.3.4). We evaluate four LLMs: GPT-4o, GPT-4o-mini(Hurst et al., 2024), GPT-3.5-Turbo (Achiam et al., 2023), and Mistral-7B-Instruct v0.3 (Jiang et al., 2023). Each model responds to each node pair over five trials, with the final estimate determined by majority voting. Full results are provided in Fig.7 in appendix, demonstrating that GPT-4o-mini and GPT-4o effectively estimate $r$, GPT-3.5-Turbo performs reasonably well, while Mistral-7B-Instruct-v0.3 fails. Balancing accuracy and cost efficiency, we select GPT-4o-mini’s estimation (Fig. 5 Right) for subsequent studies.
@@ 4.3 End-to-End Evaluation
$\bullet$ Exp.5: Main Results in the Zero-Shot Setting The main experimental results are presented in Table 2. Among the baselines, vanilla encoders and LLMs demonstrate strong zero-shot generalization. GPT-3.5-Turbo (Achiam et al., 2023) ranks first on the Sportsfit dataset, while GPT-4o (Hurst et al., 2024) achieves the best performance on Pubmed and Children.
UniGLM (Fang et al., 2024) and ZeroG (Li et al., 2024e) perform well in domains aligned with their pre-training, such as citation networks (e.g., ZeroG enhances SBert’s performance on Cora, Pubmed, and Wikics). However, both struggle on TAGs with unseen text distributions (e.g., Sportsfit) or novel graph structures (e.g., webpage networks), suggesting that fine-tuned LM encoders may suffer performance degradation on out-of-domain TAGs. Similarly, graph-SSL methods (DGI (Veličković et al., 2018), GraphMAE (Hou et al., 2022)) show limited generalization across structural shifts.
Among multi-task graph foundation models, GOFA achieves strong performance, likely benefiting from a larger pre-training corpus for graph-text alignment (Hu et al., 2021; Ding et al., 2023) compared to GraphGPT (Tang et al., 2024a) and LLaGA (Chen et al., 2024b), which are trained solely on ogbn-arxiv. However, GOFA still requires broader pre-training and instruction fine-tuning to improve generalization under text domain shifts, and its reliance on GNNs may limit effectiveness on heterophilic data.
Notably, LLM-BP and LLM-BP (appr.) achieve the highest average ranking across all datasets on both homophilic and heterophilic graphs. For fine-grained average ranking that distinguish between homophilic and heterophilic graphs, refer to Appendix. D.2. Another interesting observation is that when we randomly sample $20c$ nodes to obtain the class embeddings with the help of LLMs following Algorithm. 1, the zero-shot performance of the encoders in this setting is comparable to their performance between $5$-$10$-shot setting as shown in Table. 10 in the Appendix. Further comparisons with LLM-GNN (Chen et al., 2024e) and TEA-GLM (Wang et al., 2024) are provided in Appendix D.5.
$\bullet$ Exp.6: Main Results under Few-shot Setting We conduct the evaluation in $k=1,3,5,10$-shot settings. Using 10 different random seeds, we sample the shots from the training set and repeat the experiments 10 times. The experimental results are presented in Table 10 in Appendix D. Across all $k$-shot settings, LLM-BP and LLM-BP (appr.) outperform the baseline models.